The reading on the odometer of a car at 8:40am is 55,891km. After traveling a certain distance the odometer shows 55,912km at 8:50am. Calculate the average speed of the car in km/h
1
Expert's answer
2021-10-04T13:34:23-0400

We can find the average speed of the car as follows:


vavg=ΔsΔt,v_{avg}=\dfrac{\Delta s}{\Delta t},vavg=55912 km55891 km10 min1 h60 min=126 kmh.v_{avg}=\dfrac{55912\ km-55891\ km}{10\ min\cdot\dfrac{1\ h}{60\ min}}=126\ \dfrac{km}{h}.